# Solvable Tuple Patterns and Their Applications to Program Verification

### v1

@@ Abstract
Despite the recent progress of automated program verification techniques, fully automated verification of programs manipulating recursive data structures remains a challenge. We introduce the notion of solvable tuple patterns (STPs) to express invariants between list-like recursive data structures. A distinguishing feature of STPs is that they can be efficiently inferred from only a small number of positive samples; no negative samples are required. An SMT solver that supports the sequence theory can be used to check that an inferred STP is indeed an inductive invariant. After presenting basic properties of STPs and an STP inference algorithm, we show how to incorporate the STP inference into a CHC (Constrained Horn Clauses) solver supporting list-like data structures, which serves as a uniform backend for automated program verification tools. A CHC solver incorporating the STP inference has won the ADT-LIN category of CHC-COMP 2025 by a big margin.
@@ 1.1 Background
let rec reva l1 l2 =
  match l1 with
    [] -> l2
  | x::l1’ -> reva l1’ (x::l2)
let main l1 l2 =
  assert(reva (reva l1 l2) [] = reva l2 l1)
@@ 1.2.1 Solvable Tuple Patterns for Automated Invariant Discovery
To enable automated inference of invariant relations among list-like data structures (like the relation $l_{1}^{R}l_{2}=l_{3}$ above), we introduce the notion of solvable tuple patterns, and apply a data-driven approach to infer them. A tuple pattern is of the form $(p_{1},\ldots,p_{n})$, where $p_{i}$ consists of constants (representing elements of a sequence), variables, concatenations, the reverse operation, etc. The pattern $(p_{1},\ldots,p_{n})$ represents the set of all tuples of sequences obtained by instantiating the variables to sequences. For example, the invariant on the relation $\mathit{Reva}$ above is expressed as $(l_{1},l_{2},l_{1}^{R}l_{2})$, where $l_{1}$ and $l_{2}$ are variables representing sequences. The tuple pattern $(l_{1}l_{2},l_{2}l_{3},l_{1}l_{2}l_{3})$ expresses the set of tuples $(s_{1},s_{2},s_{3})$ where $s_{1}$ and $s_{2}$ are respectively a prefix and a postfix of $s_{3}$, and $s_{1}$ and $s_{2}$ overlap with each other in $s_{3}$ (as represented by $l_{2}$); thus (ab,bcd,abcd)(ab,bcd,abcd) [^3] Here, we denote elements by letters. belongs to $(l_{1}l_{2},l_{2}l_{3},l_{1}l_{2}l_{3})$, but $(ab,d,abcd)$ does not. The tuple patterns have been inspired from Angluin’s pattern language [2] (in fact, by using a special symbol \$\\$, a tuple pattern $(p_{1},\ldots,p_{n})$ can be expressed as a pattern $p_{1}\$\cdots\$p_{n}$). Pattern languages have a good property that they can be learned from only positive samples (in the sense of Gold [18]), but dealing with them are computationally costly: even the membership problem is NP-complete [1]. We thus introduce a subclass of tuple patterns called solvable tuple patterns. As we discuss later, the solvable tuple patterns have several pleasant properties:444 The definition of solvability and the precise meaning of the properties (i)–(iii) will be given later in Section 2. (i) like pattern languages, they can be learned from only positive samples; (ii) there is a (non-deterministic) polynomial-time algorithm for inferring a solvable tuple pattern, given positive samples; and (iii) only a small number of positive samples are required for learning a solvable tuple pattern (for example, to infer $(l_{1},l_{2},l_{1}^{R}l_{2})$, only two samples: $(ab,cd,bacd)$ and $(bc,da,cbda)$ are sufficient). Both of the patterns $(l_{1},l_{2},l_{1}^{R}l_{2})$ and $(l_{1}l_{2},l_{2}l_{3},l_{1}l_{2}l_{3})$ above belong to the class of solvable patterns.
@@
Once invariant candidates have been obtained as tuple patterns, we can use an SMT solver (that supports the sequence theory) [15, 5] to check that they are indeed inductive invariants. For the example above, to check that $\mathit{Reva}(l_{1},l_{2},l_{3})\equiv l_{3}=l_{1}^{R}l_{2}$ is indeed an inductively invariant, it suffices to check that the following formulas are valid by using an SMT solver.
$\forall l_{1},l_{2}.\;l_{2}=\epsilon^{R}l_{2}\qquad\forall x,l^{\prime}_{1},l_{2},l_{3}.\;{l^{\prime}_{1}}^{R}(x\cdot l_{2})=l_{3}\Rightarrow(x\cdot l^{\prime}_{1})^{R}l_{2}=l_{3}.$
Here, the formulas above have been obtained by substituting $l_{3}=l_{1}^{R}l_{2}$ for $\mathit{Reva}(l_{1},l_{2},l_{3})$ in the clauses (2) and (3) (and replacing primitives on lists with those on sequences). We can also confirm that the invariant is strong enough by checking the validity of the following formula, obtained from the clause (1):
$\forall l_{1},l_{2},l_{3},l_{4},l_{5}.\;l_{1}^{R}l_{2}=l_{3}\land l_{3}^{R}\epsilon=l_{4}\land l_{2}^{R}l_{1}=l_{5}\Rightarrow l_{4}=l_{5}.$
@@ 1.2.2 Combination with Arithmetic Reasoning
Whilst solvable tuple patterns have attractive features as mentioned above, invariants expressed by solvable tuple patterns are often not strong enough for the purpose of program verification. For example, consider the following program.
@@
Given an integer $n$ and a list $l$ as arguments, the function take returns the list consisting of the first (at most) $n$ elements of $l$, and drop returns the list obtained by removing the first (at most) $n$ elements from $l$. Suppose we wish to prove that $(\texttt{take}\;n\;l)@(\texttt{drop}\;n\;l)=l$ holds for every (non-negative) integer $n$ and list $l$ (where @ denotes the list append function). By using solvable tuple patterns, we can express the properties that $\texttt{take}\;n\;l$ and $\texttt{drop}\;n\;l$ respectively return a prefix and a postfix of $l$, but they are not strong enough to derive $(\texttt{take}\;n\;l)@(\texttt{drop}\;n\;l)=l$.
To address the issue above, we combine solvable tuple pattern inference with existing automated verification tools (more specifically, CHC solvers [24, 32, 9]) for reasoning about integer programs. For the example above, we construct the following integer program by abstracting lists to their lengths. [^5] As discussed in Section 3, we actually apply this kind of transformation to CHCs, rather than source programs.
@@
  if n=0 then 0 else
    if l=0 then 0 else
@@
  if n=0 then l else
    if l=0 then 0 else
@@
Here, the second argument $l$ of take’ and drop’ is an integer, representing the length of the original second argument of take and drop. We then try to prove that the abstract version of the property: $(\texttt{take'}\;n\;l)+(\texttt{drop'}\;n\;l)=l$ holds for all integers $n$ and $l$. Recent CHC solvers can quickly solve that verification problem, and can infer invariants like:
@@
$\mathit{Take}(n,l,r)\equiv\exists s.rs=l\land((\mathit{len}(l)<n\land\mathit{len}(r)=\mathit{len}(l))\lor(\mathit{len}(l)\geq n\land\mathit{len}(r)=n))$
$\mathit{Drop}(n,l,r)\equiv\exists s.sr=l\land((\mathit{len}(l)<n\land\mathit{len}(r)=0)\lor(\mathit{len}(l)\geq n\land\mathit{len}(r)=\mathit{len}(l)-n)).$
By using an SMT solver, we can confirm that these are indeed invariants strong enough to satisfy $\mathit{Take}(n,l,r_{1})\land\mathit{Drop}(n,l,r_{2})\Rightarrow r_{1}r_{2}=l$.
The above method is somewhat reminiscent of methods (such as Nelson-Oppen combination procedure [41]) for deciding a combination of theories; indeed, we are combining a method for reasoning about lists (based on solvable tuple inference) and one for verifying integer programs. Various techniques to combine static analyses have also been proposed [12, 21, 13]. To our knowledge, however, the way we combine methods for solving problems in individual domains also seems novel.
@@ 1.3 Contributions and the Organization of This Paper
  • Development of the theory and algorithm for solvable tuple patterns. We believe that the notion of solvable tuple patterns is novel and of independent interest, even outside the context of automated program verification, e.g., in the field of algorithmic learning theory.
  • Applications of solvable tuple patterns to automated verification of list-manipulating programs. As described above, we discovered a way to combine solvable tuple pattern inference with verification methods for integer-manipulating programs, which enables us to automatically verify non-trivial properties of list-manipulating programs.
@@
  • Implementation and experiments. We have implemented a new CHC solver called CHoCoL, which supports list-like data structures. We have evaluated CHoCoL by using the benchmark set of CHC-COMP 2025 ADT-LIA category, and confirmed that CHoCoL could solve many problem instances for which other CHC solvers time out. A combination of CHoCoL with another solver called Catalia [30] won the CHC-COMP 2025 ADT-LIA category by a big margin. [^6] https://chc-comp.github.io/
@@
The rest of the paper is structured as follows. In Section 2, we introduce the notion of solvable tuple patterns, present a (non-deterministic) algorithm for inferring them from positive samples, and prove that the class of solvable tuple patterns and the algorithm satisfy desired properties. Section 3 shows how to apply solvable tuple patterns to CHC solving (and hence also automated program verification). Section 4 reports an implementation and experimental results. Section 5 discusses related work and Section 6 concludes the paper.
2 Solvable Tuple Pattern Inference
In this section, we introduce the notion of solvable tuple patterns (STPs), propose an algorithm to infer STPs from positive samples, and prove properties of STPs and the algorithm. This section focuses on the learning theoretical aspect of STPs; the issues of how to collect positive samples from programs and how to apply the inferred STPs in the context of program verification will be discussed in Section 3.
We first introduce tuple patterns and give a non-deterministic inference algorithm for tuple patterns in Section 2.1. We then introduce, in Section 2.2, the class of STPs as the sound and complete characterization of the class of tuple patterns inferable by the algorithm, and show that STPs and the algorithm have nice properties. In Section 2.3, we expand the class of STPs by adding more operators (such as the reverse operator as used in the example in Section 1). We also show that we can naturally adopt STPs to express relations on sets and multisets.
2.1 Tuple Patterns and Inference Algorithm
2.1.1 Tuple Patterns.
@@ 2.1 Tuple Patterns and Inference Algorithm
Let $\Sigma$ be a set of letters, and $\mathcal{V}$ a (countably infinite) set of variables. We assume that $\Sigma$ contains at least two distinct letters. For the purpose of representing a relation among lists, $\Sigma$ is the set of possible list elements; for example, if we consider integer lists, then we let $\Sigma$ be the set of natural numbers. Below we often write $a,b,c,\ldots$ for elements of $\Sigma$.
@@
The set of tuple patterns, ranged over by $t$, is defined by:
$t::=(p_{1},\ldots,p_{k})\qquad p\in(\Sigma\cup\mathcal{V})^{*}.$
@@
A tuple pattern $t=(p_{1},\ldots,p_{k})$ represents a set of $k$-tuples of $\Sigma$-sequences, i.e., a $k$-ary relation between sequences, obtained by replacing each variable with a sequence of letters in $\Sigma$. For example, $(x,y,xy)$ represents a set of triples of the form $(s_{1},s_{2},s_{1}s_{2})$ where $s_{1},s_{2}\in\Sigma^{*}$. Formally, we define $\mathcal{L}(t)\subseteq\Sigma^{*}\times\cdots\times\Sigma^{*}$ as:
$\{[s_{1}/x_{1},\ldots,s_{n}/x_{n}]t\mid s_{1},\ldots,s_{n}\in\Sigma^{*}\}$
where $x_{1},\ldots,x_{n}$ are the variables occurring in $t$, and $s_{1}/x_{1},\ldots,s_{n}/x_{n}]$ denotes the simultaneous substitution of $s_{1},\ldots,s_{n}$ for $x_{1},\ldots,x_{n}$.
@@
We identify tuple patterns up to variable renaming. For example, we identify $(xy,x)$ with $(yx,y)$. We write $\mathbf{Vars}(p)$ for the set of variables occurring in $p$. We write $p_{1}\cdot p_{2}$ for the concatenation of two patterns sequences $p_{1}$ and $p_{2}$, and often omit $\cdot$, writing $p_{1}p_{2}$. We write $|p|$ for the length of a pattern $p$. For a tuple pattern $t=(p_{1},\ldots,p_{k})$, we write $|t|$ and #t\#{t} respectively for $k$ and $|p_{1}\cdots p_{k}|+k$.
@@
We will later (in Section 2.3) extend tuple patterns with other pattern constructors such as $\cdot^{R}$, to accommodate the $\mathit{Reva}$ example in Section 1.
@@
In the case of a singleton tuple pattern $(p)$, our definition of $\mathcal{L}$ coincides with the definition of a variation of pattern languages [2, 1] called E-pattern languages (or erasing pattern languages) [49]. In the definition of ordinary pattern languages [2, 1], $\mathcal{L}(p)$ is defined as $\{[s_{1}/x_{1},\ldots,s_{n}/x_{n}]p\mid s_{1},\ldots,s_{n}\in\Sigma^{+}\}$; empty strings cannot be substituted for variables. In view of the application to program verification discussed in Section 3, we allow substitutions of empty strings, which correspond to empty lists. ∎
Notation 2.1.
We write $\widetilde{\cdot}$ for a sequence. By abuse of notation, we write $\widetilde{p}$ for a sequence $p_{1},\ldots,p_{n}$ and also for a tuple pattern $(p_{1},\ldots,p_{n})$, depending on the context; similarly for sequences $\widetilde{x}$ and $\widetilde{s}$ of variables and strings respectively.
2.1.2 Tuple Pattern Inference Problem.
Learning data, denoted by $M$, is an $m\times n$-matrix consisting of elements of $\Sigma^{*}$. We write $M[i][j]$ for the element in the $i$-th row and $j$-th column. We also write $M[i]$ for $(M[i][1],\ldots,M[i][n])$, and M[][j]M[*][j] for $(M[1][j],\ldots,M[m][j])$. We write $\mathit{shape}(M)$ for $(m,n)$, $\mathit{rows}({M})$ for $m$, and $\mathit{size}(M)$ for $\sum_{i\in\{1,\ldots,m\},j\in\{1,\ldots,n\}}(1+|M[i][j]|)$, where $|s|$ denotes the length of string $s\in\Sigma^{*}$.

### v2

@@ Abstract
Despite the recent progress of automated program verification techniques, fully automated verification of programs manipulating recursive data structures remains a challenge. We introduce solvable tuple patterns (STPs) and conjunctive STPs (CSTPs), novel formalisms for expressing and inferring invariants between list-like recursive data structures. A distinguishing feature of STPs is that they can be efficiently inferred from only a small number of positive samples; no negative samples are required. After presenting properties and inference algorithms of STPs and CSTPs, we show how to incorporate the CSTP inference into a CHC (Constrained Horn Clauses) solver supporting list-like data structures, which serves as a uniform backend for automated program verification tools. A CHC solver incorporating the (C)STP inference has won the ADT-LIN category of CHC-COMP 2025 by a significant margin.
@@ 1.1 Background
let rec reva l1 l2 = match l1 with [] -> l2 | x::l1’ -> reva l1’ (x::l2)
let main l1 l2 = assert(reva (reva l1 l2) [] = reva l2 l1)
@@ 1.2.1 Solvable Tuple Patterns for Automated Invariant Discovery
[…]
@@
To apply STPs to automated verification of list-manipulating programs, we consider CHCs on words, which are Horn clauses extended with word equality/inequality constraints. The clauses on $\mathit{Reva}$ above may be considered CHCs on words, by viewing lists as words. We show (i) given a set of CHCs on words consisting of definite clauses, there exists an algorithm to compute the least model describable by conjunctive STPs (i.e., a conjunction of a finite number of STPs), [^5] Stated without using the terminology of CHCs, this means that there exists an algorithm to compute the strongest inductive invariant among those expressible by using CSTPs, given a list-manipulating program with loops and first-order recursive functions. and (ii) it is decidable whether, given a system of CHCs over words, it has a model describable by conjunctive STPs. We can directly apply these results to the $\mathit{Reva}$ example, and automatically obtain the invariant: $\mathit{Reva}(l_{1},l_{2},l_{3})\equiv l_{3}=l_{1}^{R}l_{2}$.
@@ 1.2.2 Combination with Arithmetic Reasoning
Whilst CSTPs have attractive features as mentioned above, invariants expressed by CSTPs are not always strong enough for the purpose of program verification, especially for programs using a combination of lists and integers. For example, consider the following program.
@@
Given an integer $n$ and a list $l$ as arguments, the function take returns the list consisting of the first (at most) $n$ elements of $l$, and drop returns the list obtained by removing the first (at most) $n$ elements from $l$. Suppose we wish to prove that $(\texttt{take}\;n\;l)@(\texttt{drop}\;n\;l)=l$ holds for every (non-negative) integer $n$ and list $l$ (where @ denotes the list append function). By using solvable tuple patterns, we can infer the properties that $\texttt{take}\;n\;l$ and $\texttt{drop}\;n\;l$ respectively return a prefix and a suffix of $l$, but they are not strong enough to derive $(\texttt{take}\;n\;l)@(\texttt{drop}\;n\;l)=l$.
To address the issue above, we combine STP inference with existing automated verification tools (more specifically, CHC solvers over integer arithmetic [25, 35, 8]). For the example above, we construct the following integer program by abstracting lists to their lengths. [^6] As discussed in Section 5, we actually apply this kind of transformation to CHCs, rather than source programs.
@@
  if n=0 then 0 else if l=0 then 0 else
@@
  if n=0 then l else if l=0 then 0 else
@@
Here, the second argument $l$ of take’ and drop’ is an integer, representing the length of the second argument of the original functions take and drop. We then try to prove that the abstract version of the property: $(\texttt{take'}\;n\;l)+(\texttt{drop'}\;n\;l)=l$ holds for all integers $n$ and $l$. State-of-the-art CHC solvers can quickly solve that verification problem, and can infer invariants like:
@@
$\mathit{Take}(n,l,r)\equiv\exists s.rs=l\land((|l|<n\land|r|=|l|)\lor(|l|\geq n\land|r|=n))$
$\mathit{Drop}(n,l,r)\equiv\exists s.sr=l\land((|l|<n\land|r|=0)\lor(|l|\geq n\land|r|=|l|-n)).$
These are indeed invariants strong enough to satisfy $\mathit{Take}(n,l,r_{1})\land\mathit{Drop}(n,l,r_{2})\Rightarrow r_{1}r_{2}=l$.
The above method is somewhat reminiscent of methods (such as Nelson-Oppen combination procedure [46]) for deciding combinations of theories; indeed, it combines a method for reasoning about lists (based on STP inference) with one for verifying integer programs. Various techniques to combine static analyses have also been proposed [11, 22, 12]. To our knowledge, however, the way we combine methods for solving problems in individual domains also seems novel.
@@ 1.3 Contributions and the Organization of This Paper
  • Development of the theory and algorithms for solvable tuple patterns (STPs) and conjunctive STPs (CSTPs). We believe that the notions of STPs and CSTPs are novel and of independent interest, even outside the context of automated program verification, e.g., in the field of algorithmic learning theory.
  • Applications of solvable tuple patterns to automated verification of list-manipulating programs. To this end, we introduce CHCs over words, and show that it is decidable whether, given a system of CHCs over words, it has a model describable by CSTPs. We also consider CHCs over multisets and sets, and show that it is decidable whether a given system of CHCs over multisets or sets has a model describable by the multiset or set version of CSTPs. We also propose a method to strengthen this approach by combining it with verification methods for integer-manipulating programs.
@@
  • Implementation and experiments. Based on the results above, we have implemented a new CHC solver called CHoCoL, which supports list-like data structures. We have evaluated CHoCoL by using the benchmark set of the CHC-COMP 2025 ADT-LIA category, and confirmed that CHoCoL significantly outperformed other CHC solvers. A combination of CHoCoL with another solver called Catalia [33] won the CHC-COMP 2025 ADT-LIA category by a large margin. [^7] https://chc-comp.github.io/
@@
The rest of the paper is structured as follows. In Sections 24, we introduce the notion of solvable tuple patterns (STPs) and conjunctive STPs (CSTPs), present algorithms for inferrig them from positive samples, and prove that the classes of STPs and CSTPs satisfy desired properties. Section 5 shows how to apply CSTPs to CHC solving (and hence also automated program verification). Section 6 reports an implementation and experimental results. Section 7 discusses related work and Section 8 concludes the paper.
2 Tuple Patterns and Inference Algorithm
In this section, we introduce the notions of tuple patterns and conjunctive tuple patterns, and provide inference algorithms. In the next section, we will introduce a subclass of tuple patterns called solvable tuple patterns as the sound and complete characterization of the class of tuple patterns inferable by the algorithm.
2.1 Tuple Patterns and Conjunctive Tuple Patterns
@@
Let $\Sigma$ be a set of letters, and $\mathcal{V}$ a (countably infinite) set of variables. We assume that $\Sigma$ contains at least two distinct letters. For the purpose of representing a relation among lists, $\Sigma$ is the set of possible list elements; for example, if we consider integer lists, then we let $\Sigma$ be the set of integers. Below we often write $a,b,c,\ldots$ for elements of $\Sigma$. We write $\mathit{dom}(f)$ for the domain of a map $f$.
@@
The sets of tuple patterns and conjunctive tuple patterns, ranged over by $t$ and $\gamma$ respectively, are defined by:
$t\mbox{ (tuple patterns) }::=(p_{1},\ldots,p_{k})\qquad p\in(\Sigma\cup\mathcal{V})^{*}$
$\gamma\mbox{ (conjunctive tuple patterns) }::=t_{1}\land\cdots\land t_{m}.$
@@
A tuple pattern $t=(p_{1},\ldots,p_{k})$ represents a set of $k$-tuples of $\Sigma$-sequences, i.e., a $k$-ary relation between sequences, obtained by replacing each variable with a sequence of letters in $\Sigma$. For example, $(x,y,xy)$ represents the set of triples of the form $(s_{1},s_{2},s_{1}s_{2})$ where $s_{1},s_{2}\in\Sigma^{*}$. Formally, we define $\mathcal{L}(t)\subseteq\Sigma^{*}\times\cdots\times\Sigma^{*}$ as: $\{[s_{1}/x_{1},\ldots,s_{n}/x_{n}]t\mid s_{1},\ldots,s_{n}\in\Sigma^{*}\}$ where $x_{1},\ldots,x_{n}$ are the variables occurring in $t$, and $s_{1}/x_{1},\ldots,s_{n}/x_{n}]$ denotes the simultaneous substitution of $s_{1},\ldots,s_{n}$ for $x_{1},\ldots,x_{n}$. A conjunctive tuple pattern $t_{1}\land\cdots\land t_{m}$ (where $m\geq 1$) represents the intersection of the sets of tuples represented by $t_{1},\ldots,t_{m}$, as defined by: $\mathcal{L}(t_{1}\land\cdots\land t_{m})=\mathcal{L}(t_{1})\cap\cdots\cap\mathcal{L}(t_{m})$. For example, $(x,y,xz)\land(x,y,yz)$ represents the set of triples $(s_{1},s_{2},s_{3})$ where both $s_{1}$ and $s_{2}$ are prefixes of $s_{3}$; here, note that variables are implicitly bound in each conjunct: the two occurrences of $z$ in $(x,y,xz)$ and $(x,y,yz)$ may be distinct. We write $\widetilde{\cdot}$ for a sequence. By abuse of notation, we write $\widetilde{p}$ for a sequence $p_{1},\ldots,p_{n}$ and also for a tuple pattern $(p_{1},\ldots,p_{n})$, depending on the context; similarly for sequences $\widetilde{x}$ and $\widetilde{s}$ of variables and strings respectively.
@@
We identify tuple patterns up to variable renaming. For example, we identify $(xy,x)$ with $(yx,y)$. We write $\mathbf{Vars}(p)$ and $\mathbf{Vars}(t)$ for the sets of variables occurring in $p$ and $t$ respectively. We write $p_{1}\cdot p_{2}$ for the concatenation of two patterns sequences $p_{1}$ and $p_{2}$, and often omit $\cdot$, writing $p_{1}p_{2}$. We write $|p|$ for the length of a pattern $p$. For a tuple pattern $t=(p_{1},\ldots,p_{k})$, we write $|t|$ and #t\#{t} respectively for $k$ and $|p_{1}\cdots p_{k}|+k$.
@@
We will later (in Section 4) extend tuple patterns with the pattern $x^{R}$ (which represents the reverse of $x$), to accommodate the $\mathit{Reva}$ example in Section 1.
@@
In the case of a singleton tuple pattern $(p)$, our definition of $\mathcal{L}$ coincides with the definition of a variation of pattern languages [2, 1] called E-pattern languages (or erasing pattern languages) [57]. In the definition of original pattern languages [2, 1] (called NE-pattern languages), $\mathcal{L}(p)=\{[s_{1}/x_{1},\ldots,s_{n}/x_{n}]p\mid s_{1},\ldots,s_{n}\in\Sigma^{+}\}$; empty strings cannot be substituted for variables. In view of the application to program verification discussed in Section 5, we allow substitutions of empty strings, which correspond to empty lists. ∎
2.2 Tuple Pattern Inference Problem
Learning data, denoted by $M$, is an $m\times n$-matrix consisting of elements of $\Sigma^{*}$. We write $M[i][j]$ for the element in the $i$-th row and $j$-th column. We also write $M[i]$ for $(M[i][1],\ldots,M[i][n])$, and M[][j]M[*][j] for $(M[1][j],\ldots,M[m][j])$. We write $\mathit{shape}(M)$ for $(m,n)$, $\mathit{rows}({M})$ for $m$, $\mathit{dim}({M})$ for $n$, and $\mathit{size}(M)$ for $\sum_{i\in\{1,\ldots,m\},j\in\{1,\ldots,n\}}(1+|M[i][j]|)$, where $|s|$ denotes the length of string $s\in\Sigma^{*}$. We sometimes regard learning data $M$ as the set $\{M[1],\ldots,M[m]\}$, and write $M_{1}\subseteq M_{2}$ when $M_{1}$ is a subset of $M_{2}$ as sets.